If 3 gallons of Grade A oil are mixed with 7 gal of grade B oil the resulting mixture is 43 cents/gal.However,if 3 gal of Grade A oil are mixed with 2 gal of Grade B oil the resulting mixture is worth 46 cent/gal.Find the price per gallon of each grade.
Thanks for the help.!

Answer by richwmiller(17219)

3a+7b=43*10,
3a+2b=46*5
a = 50, b = 40